A motorboat leaves Galveston, TX. One-half hour after the boat leaves, a second boat leaves Galveston following the first boat. The first boat is traveling at 25 mph, while the second boat travels at 35 mph. How long will it take the second boat to over take the first boat?
Answer by rwm(914)

basically the first boat got ahead by 12.5 miles
the second boat gains on the first boat at 10 mph.
how long will it take the second boat to go 12.5 miles gaining at 10 mph
r*t=d
25*(t+1/2)=35t
25t+12.5=35t
12.5=10t
1.25=t
1 hour 15 minutes